OpenStudy (anonymous):

-1.2=0.9x-3.2x+3.4 -1.2=-2.3x+3.4 Is this part right? Now what?

OpenStudy (dumbcow):

subtract 3.4 on both sides

OpenStudy (dumbcow):

then divide by -2.3

OpenStudy (anonymous):

Ok. So x=-2, right?

OpenStudy (dumbcow):

no you should get -4.6 = -2.3x then you are diving a neg by a neg -> positive x = 2

OpenStudy (anonymous):

x = 2

OpenStudy (anonymous):

Ok. Thank you.

OpenStudy (anonymous):

-1.2= 0.9x - 3.2x + 3.4 -1.2= -2.3x + 3.4 2.3x = 3.4 + 1.2 = 4.6 x = 4.6 = 2.3 = 2

OpenStudy (anonymous):

i mean x = 4.6 / 2.3 = 2